# Neural Transparency: Mechanistic Interpretability Interfaces for Anticipating Model Behaviors for Personalized AI

### v1

@@ 5.2. The Transparency Paradox: High Value to Users, Null Behavioral Effects
The central paradox in our findings is that the persona visualization was highly valued by users yet produced no measurable effects on their design behaviors or outcomes. Participants in the visualization condition rated the tool as highly helpful and expressed strong desire to use it again, yet showed no significant differences from controls in several metrics: number of prompt iterations, magnitude of persona changes, message engagement, or confidence in predicting chatbot behaviors.
This disconnect between perceived value and behavioral impact suggests several possibilities. First, the visualization may have influenced aspects of the design process we did not measure. Our metrics focused on quantifiable behaviors (iteration counts, trait changes, message volumes), but the visualization may have primarily affected the quality of users’ mental models or their subjective experience of agency and control. The qualitative data supports this interpretation: visualization group users described a more intentional design process compared to control users, even though both groups ultimately achieved similar satisfaction levels.
@@
Second, our experimental task may not have been challenging enough to reveal the visualization’s benefits. Participants were creating emotional support chatbots, a relatively benign use case where most reasonable approaches would likely produce acceptable results. The safety-relevant traits we measured (toxicity, sycophancy, hallucination) may not have been salient enough in this context to drive design changes. Future work with more adversarial tasks (e.g., recreating problematic personas or designing chatbots for contexts where specific traits are critical) might reveal stronger behavioral effects.
@@
Third, the visualization may require iterative exposure to influence behavior. Our single-session design gave users limited opportunity to internalize the relationship between system prompts and predictions, build expertise with the visualization, or develop strategies for using it effectively. Longitudinal studies tracking users across multiple chatbot creation sessions could reveal whether the visualization’s impact grows with experience.
@@
Finally, we could have implemented cognitive forcing functions — an intervention to disrupt heuristic, automatic, thinking — to have participants carefully reason about the consequences of their system prompt on the persona visualization. For example, we could have used tooltips to remind users to think critically at what language in their system prompt elicited persona scores. Future work may also explore highlighting how specific words and phrases had influence over the system prompt, creating causal inks between language and behavior.
5.3. Trust Through Transparency: Mechanisms and Implications
Despite null effects on design behavior, the visualization significantly increased post-interaction trust. This trust increase is particularly noteworthy because it occurred after users had already interacted with their chatbot for 10 minutes—a period during which they could directly observe whether the predictions matched actual behavior. The increase in trust was not mediated by increased confidence in predicting behaviors (which showed no group differences), suggesting the visualization increased trust through transparency itself rather than through improved behavioral understanding.
Qualitative analysis revealed how this trust developed through fundamentally different design processes. Control participants described discovery-oriented processes marked by surprise and reactive problem-solving, treating unexpected behaviors as system failures requiring correction. Visualization participants described intentional processes characterized by proactive adjustment, treating prediction-behavior mismatches as learning opportunities. One visualization user reflected, “Changing the prompt completely changed the character more than I anticipated…I should’ve probably been more clear as to what I wanted on the spectrum”, internalizing lessons about prompt specificity rather than blaming the system.
Both groups wanted more time, but for different reasons: control participants needed time to discover and correct unexpected behaviors, while visualization participants wanted to explore and refine already-satisfactory designs—reactive correction versus proactive exploration.
This finding has important practical implications. If transparency tools can increase user trust without requiring changes to design workflows or imposing additional cognitive burden, they become more viable for real-world deployment. However, this raises normative questions about whether increased trust without improved behavioral prediction is desirable. This concern is partially mitigated by our finding that visualization users’ qualitative reflections showed more sophisticated understanding of prompt-behavior relationships and more metacognitive awareness about calibration challenges, even if quantitative metrics failed to capture these process differences.
@@ 5.4. Limitations and Constraints
Several limitations constrain generalizability of our findings. First, we relied on GPT-4.1-mini to evaluate trait expression in responses from Llama-3.2-3B-Instruct, introducing potential biases from LLM limitations in capturing the nuances of trait expression in language. While this automated evaluation enabled the scale necessary for persona vector generation, it may not fully reflect human judgments of trait expression.
@@
Third, we generated persona vectors from model responses rather than from system prompts directly. While our linear regression analysis demonstrate these representations can evaluate traits in system prompts, this task difference may introduce inaccuracy. Investigating whether generating persona vectors directly from system prompt activations improves efficacy would strengthen the methodology.
Fourth, our 10-minute interaction period, while sufficient to gather initial impressions, may not capture longer-term dynamics of chatbot relationships. Cases of problematic AI companion interactions typically emerge over weeks or months of use, suggesting that single-session studies may systematically underestimate safety risks and the value of transparency tools.
Fifth, participants’ inaccurate estimations of trait activations may have been due to transforming coarse ordinal data (activations from 0-5) to normalized values between 0-1. This may have systematically shifted participants’ predictions to be more extreme than they otherwise would have been if they were offered more ordinal granularity for prediction. Future research should explore if people are actually more calibrated than this work suggests by utilizing a survey that probes each pole of the persona vector individually with more granularity (e.g. choices 0-10 for each polar-end of the persona vectors).
Nonetheless, we believe this work shows promise for translating neural-level understanding into practical tools for everyday users.
5.5. Future Directions: Toward Interpretability-Informed Design
@@ 5.5.2. Adversarial task conditions:
Design experimental tasks where transparency tools become necessary rather than optional—contexts where users must avoid specific dangerous traits, satisfy strict behavioral requirements, or debug problematic personas. Such studies could reveal when transparency crosses the threshold from “nice to have” to “essential for success.”
5.5.3. Comparative transparency mechanisms:
Systematically compare persona vectors to other interpretability techniques (linear probes, sparse autoencoders, circuit analysis) for user-facing applications. Different methods may offer different tradeoffs between accuracy, computational cost, and interpretability.
@@ 5.5.4. Active steering interfaces:
Extend beyond passive prediction to active control, allowing users to directly manipulate persona vector activations to steer behavior. While prior work shows capability degradation with extreme steering values (Chen et al., 2025), users might prefer direct control despite these limitations.
5.5.5. Standardized disclosure frameworks:
Develop “AI nutritional labels” or standardized trait disclosure systems that could be adopted across chatbot platforms. If trait/behavioral disclosures became normalized parts of AI interfaces, users could develop literacy in interpreting them, potentially improving calibration over time.
@@ 5.5.6. Vulnerable population studies:
5.6. Broader Implications: Interpretability for the End User
This work represents a step toward democratizing mechanistic interpretability—translating techniques developed by and for AI researchers into tools accessible to end users. The enthusiastic reception of our visualization suggests a desire for neural-level transparency even among non-technical users, challenging assumptions that such information must be hidden to avoid overwhelming or confusing users.
Our findings reveal a fundamental challenge in translating transparency into behavioral impact. While participants could see inside the model through our visualization, this visibility alone didn’t translate into more effective control. This apparent limitation, however, may reflect our experimental design rather than the visualization itself. Emotional support chatbots represent a convergent design space where most reasonable approaches achieve similar outcomes—empathetic, supportive, and non-toxic personalities. In such constrained tasks, transparency tools have limited room to demonstrate behavioral impact because the “correct” design choices are already apparent.
The trust findings raise important questions about the normative goals of transparency. Should interpretability tools primarily aim to improve user performance (helping them create better chatbots), or is transparency valuable in itself for supporting user autonomy and informed consent? Our results suggest these goals may sometimes diverge: users valued transparency even when it didn’t improve their behavioral predictions. This suggests interpretability might serve values beyond instrumental task performance–the ability to understand the complex and previously opaque systems that influence our lives.

### v2

@@ 5.2. High Value to Users, Null Behavioral Effects
A contradiction in our findings is that the persona visualization was valued by users yet produced no measurable effects on their design behaviors or outcomes. Participants in the visualization condition consistently rated the tool as helpful and expressed strong desire to use it again, yet showed no significant differences from controls in several metrics: number of prompt iterations, between-group persona score differences, message engagement, or confidence in predicting chatbot behaviors.
@@
This disconnect between perceived value and behavioral impact suggests several possibilities. To begin with, our experimental task may not have been challenging enough to reveal the visualization’s benefits. Participants were creating emotional support chatbots, a relatively straightforward use case where most reasonable approaches would likely produce acceptable results. The safety-relevant traits we measured (toxicity, sycophancy, hallucination) may not have been salient enough in this context to drive design changes. Future work with more adversarial tasks (e.g., recreating problematic personas or designing chatbots for contexts where specific traits are critical) might reveal stronger behavioral effects from the visualization.
@@
Furthermore, the visualization may require iterative exposure to influence behavior. Our single-session design gave users limited opportunity to internalize the relationship between system prompts and predictions, build expertise with the visualization, or develop strategies for using it effectively. Longitudinal studies tracking users across multiple chatbot creation sessions could reveal whether the visualization’s impact grows with experience.
@@
Finally, the visualization may have benefited from the addition of cognitive forcing functions—an intervention to disrupt heuristic, automatic, thinking—so that participants carefully reason about the consequences of their system prompt on the persona visualization. One could implement such a design using tooltips to remind users to think critically. Future work may also explore highlighting how specific words and phrases had influence over the system prompt, creating causal inks between language and behavior.
@@ 5.3. Limitations and Constraints
Several features of our experiment warrant further exploration in order to improve the generalizability of our findings. First, we relied on GPT-4.1-mini to evaluate trait expression in responses from Llama-3.2-3B-Instruct, introducing potential limitations in capturing the nuances of trait expression in language. While this automated evaluation enabled the scale necessary for persona vector generation, it may not fully reflect human judgments of trait expression.
@@
Third, our 10-minute interaction period, while sufficient to gather initial impressions, may not capture longer-term dynamics of chatbot relationships. Cases of problematic AI companion interactions typically emerge over weeks or months of use, suggesting that single-session studies may systematically underestimate safety risks and the value of transparency tools.
Fourth, participants’ inaccurate estimations of trait activations may have been due to transforming coarse ordinal data (activations from 0-5) to normalized values between 0-1. This may have systematically shifted participants’ predictions to be more extreme than they otherwise would have been if they were offered more ordinal granularity for prediction. Future research should explore if people are actually more calibrated than this work suggests by utilizing a survey that probes each pole of the persona vector individually with more granularity (e.g. choices 0-10 for each polar-end of the persona vectors).
5.4. Future Directions
@@ 5.4.2. Adversarial task conditions:
Design experimental tasks where transparency tools become necessary so users must avoid specific dangerous traits, satisfy strict behavioral requirements, or debug problematic personas. Such studies could reveal when transparency goes from a convenience to vital for the task.
@@ 5.4.3. Active steering interfaces:
Extend beyond passive prediction to active control, allowing users to directly manipulate persona vector activations to steer behavior (allbert2024identifying). While prior work shows capability degradation with extreme steering values (chen2025persona), users might prefer direct control despite these limitations.
5.4.4. Standardized disclosure framework:
Develop a standardized trait disclosure framework that could be adopted across chatbot platforms. If trait/behavioral disclosures became normalized parts of AI interfaces, users could develop literacy in interpreting them, potentially improving calibration over time. This may be thought of as a ‘nutritional label’ for chatbots that can be referenced during interaction.
@@ 5.4.5. Vulnerable population studies:
5.5. Broader Implications
This work represents a step toward operationalizing mechanistic interpretability—translating techniques developed by and for AI researchers into tools accessible to end users. The enthusiastic reception of our visualization suggests a desire for neural-level transparency even among non-technical users, challenging assumptions that such information must be hidden to avoid overwhelming or confusing users.
Our findings related to trust after interaction suggest an important distinction between trust and control. Participants trusted the system more after using the visualization, yet showed no corresponding increase in confidence about predicting its behavior. Rather than causing overconfidence, the visualization seems to build trust through procedural transparency—users saw measurable evidence of how system prompts translate to neural activations. This suggests the visualization helped users trust the system rather than inflating their sense of control. Predictions from interpretable mechanisms may enable more informed decisions about AI use even when behaviors remain complex and difficult to perfectly anticipate. Therefore, transparency can be valuable not because it guarantees control, but because it supports informed consent.